# RA-DP: Rapid Adaptive Diffusion Policy for Training-Free High-frequency Robotics Replanning

### v1

@@ Abstract
Diffusion models exhibit impressive scalability in robotic task learning, yet they struggle to adapt to novel, highly dynamic environments. This limitation primarily stems from their constrained replanning ability: they either operate at a low frequency due to a time-consuming iterative sampling process, or are unable to adapt to unforeseen feedback in case of rapid replanning. To address these challenges, we propose RA-DP, a novel diffusion policy framework with training-free high-frequency replanning ability that solves the above limitations in adapting to unforeseen dynamic environments. Specifically, our method integrates guidance signals which are often easily obtained in the new environment during the diffusion sampling process, and utilizes a novel action queue mechanism to generate replanned actions at every denoising step without retraining, thus forming a complete training-free framework for robot motion adaptation in unseen environments. Extensive evaluations have been conducted in both well-recognized simulation benchmarks and real robot tasks. Results show that RA-DP outperforms the state-of-the-art diffusion-based methods in terms of replanning frequency and success rate. Moreover, we show that our framework is theoretically compatible with any training-free guidance signal.
@@ I Introduction
Diffusion models have gained widespread attention for their effective performance in various tasks, such as image generation [1], video synthesis [2], and text-to-image conversion [3]. The ability of diffusion models to learn complex data distributions has made them a powerful tool in creative and predictive applications. More recently, diffusion models have been employed for learning and improving robotic control tasks [4, 5, 6, 7]. In this context, they have demonstrated the expressiveness needed for complex manipulation tasks, stable training, and adaption to higher-dimensional action spaces. Consequently, these attributes make diffusion models ideal for addressing a wide range of challenges in robotic control.
@@
Figure 1: Comparison between DP and RA-DP. DP exhibits limited replanning frequency and fails to adapt to unforeseen dynamic observations, whereas RA-DP replans actions using the latest dynamic feedback at every diffusion denoising step in a training-free manner. The lighter the shade of blue, the lower the level of Gaussian noise. A white action represents an executable action.
For robot control, the authors of [4] propose Diffusion Policy (DP), which typically generates 16 actions by iterative sampling, executes the first 8 actions, and then replans conditioned on the new observations, resulting in a replanning frequency of approximately $37$ Hz, which is substantially slower compared to standard low-level robotic control systems. As a result, DP can create a bottleneck for the entire system. One way to increase the replanning rate is through shortening prediction and action horizons at the cost of lowering the success rate. Alternatively, one can resort to fast diffusion sampling techniques, such as consistency models [7], to improve frequency. However, these approaches still struggle to adapt to novel environment feedback, such as unforeseen obstacles.
@@
RA-DP establishes a fixed-length action queue. During training, each action within the queue is independently perturbed with varying noise levels, unlike previous methods that use uniform noise levels across actions. A standard U-Net is trained to predict clean actions conditioned on slowly changing feedback (e.g., a fixed goal position). During inference, RA-DP maintains an action queue with monotonically increasing noise levels. Other than conditioning on the foreseen, slow-changing feedback during the training stage, each action is also seamlessly steered by the training-free guidance from possible rapid changing environment feedback. RA-DP removes the executable clean action from the front (dequeue) of the queue and appends a new noisy one to the tail (enqueue) at every denoising step, thereby achieving high-frequency training-free replanning. For example, in goal-reaching tasks in cluttered environments, RA-DP is able to incorporate live obstacle data to dynamically guide the actions.
@@
  • We propose RA-DP, a novel diffusion-based framework for robotic manipulation with a training-free high-frequency replanning to adapt to fast changing environments.
@@ II-A Diffusion Models in Robotics
While these diffusion-based approaches excel in static or slowly changing environments, they fall short in rapidly evolving real-world scenarios. This is primarily due to their limited replanning frequency, hindering their application to tasks, such as collision avoidance in dynamic settings. To address this issue, we propose a novel high-frequency replanning diffusion-based framework that benefits from a training-free loss-based guidance in order to enhance adaptability and efficiency of the controller in highly dynamic environments.
@@ II-C Fast Replanning in Diffusion-based Controllers
Diffusion models [16, 17] generate high-quality samples through iterative refinement from Gaussian noise. However, this iterative process leads to low replanning frequencies, limiting their real-time performance. To overcome this challenge, the approaches, such as DDIM [18], DPM-Solver [19], and Genie [20] use advanced ordinary differential equation (ODE) solvers to reduce the number of sampling steps needed for high-quality samples. Despite such an improvement, these methods still fail to achieve the high-frequency replanning required for low-latency decision-making applications.
More recent consistency policy [7] leverages diffusion model distillation to enable single-step sampling via direct noise-to-data mappings. Although this approach facilitates fast replanning, it has a number of shortcomings. The additional distillation stage applied to pretrained diffusion policies increases computational cost; The single-step sampling process compromises accuracy compared to iterative approaches; Lastly, the single-step architecture lacks intermediate latent representations, precluding the application of training-free, loss-based guidance techniques [10] necessary for adapting to unforeseen conditions. This is due to the fact that applying guidance directly to the data space either causes divergence or fails to influence the sampling process effectively.
In this work, we propose a significantly different approach, in which we integrate the action queue with different noise levels for high-frequency replanning. More specifically, our training and inference pipeline denoises the action queue with monotonically increasing noise levels, enabling single-step action generation without compromising sample quality. Additionally, we retain intermediate latent representations for training-free, loss-based guidance, thereby supporting adaption to novel scenarios during inference. These characteristics make our proposed method compatible with any existing fast diffusion solvers.
@@
Similarly, hierarchical policies, such as Yell At Your Robot (YAY) [22], utilize high-level human language interventions as real-time corrective feedback for low-level policy errors. However, YAY’s VAE-based language-conditioned behavior cloning policy exhibits constrained action generation capabilities compared with diffusion-based approaches. Furthermore, while YAY requires continuous training to integrate language feedback, our method employs a training-free guidance mechanism.
@@ III Proposed Method
We begin by introducing the notation used throughout this paper and review key aspects of Diffusion Policy (DP) in Section III-A. Then, in Section III-B, we present the training method of our RA-DP framework, which manages actions with independent noise levels during forward diffusion processes as opposed to how traditional DP operates. Lastly, in Section III-C, we provide a detailed explanation of the inference process of our framework and demonstrate how it effectively achieves adaptive high-frequency replanning.
@@ III-A Notation and Diffusion Policy (DP)
We emphasize that the standard DDPM training used in diffusion treats the entire prediction sequence $A_{t}$ as a single example, where each action $a_{t}$ in the sequence is perturbed equally with the same noise level during the forward diffusion process. Instead, if we were to sample a sequence of actions during inference with varying noise level, the policy trained with standard DDPM diverges and fails to predict normal actions. This is due to the mismatch between training and inference paradigm. However, we argue that varying noise level is needed to achieve real-time close-loop control.
@@ III-C Adaptive High-frequency Replanning Process
Finally, we update $\hat{A}_{t}^{\mathbf{k-1}}$ by appending a new noisy action $\hat{a}_{t+H}^{H\tau}$ to get the new action queue $\hat{A}_{t+1}^{\mathbf{k}}$. Leveraging our proposed action queue sampling method, at each time step $t$ after just one denoising step, we obtain an noiseless excutable action $\hat{a}_{t}^{0}$ conditioned on the most recent observation $O_{t}$.
High-Frequency Replanning. Let $\Delta_{p}$ denote the cost of a one-step DDIM sampling and $\Delta_{a}$ the cost of a one-step action execution. Then, the replanning frequency of our framework is given by $\frac{1}{\Delta_{p}+\Delta_{a}}$. Notably, it preserves sampling quality unlike other single-step diffusion samplers, as the total number of sampling steps for each action remains $H$.
@@
Adaption to Novel environment Feedback. Beyond conditioning on the slow-changing observation $O_{t}$, genuine high-frequency replanning frameworks must swiftly respond to unforeseen environmental feedback. To this end, we propose to integrate a training-free loss-based guidance module [25, 26, 9], which outputs guidance gradients from arbitrary predefined constraints that influences the diffusion sampling process, thus achieving adaption to novel environment feedback (see Fig.2). Moreover, the guidance module can be seamlessly integrated with the aforementioned action queue mechanism.
We describe the guided diffusion model from the perspective of score matching-based generative model, of which both the forward and reverse diffusion processes are formulated as stochastic differential equations (SDEs), in which train the U-Net to parameterize the score function $\nabla_{A^{\mathbf{k}}}\log p_{\theta}(A^{\mathbf{k}})$. However, a guided sampling process aims to sample $A^{\mathbf{k}}$ conditioned on the guidance signal $\mathcal{G}$, i.e., it requires the score function:
@@ IV Experiments
We evaluate RA-DP across simulated and real-robot scenarios to highlight the importance of high-frequency replanning and demonstrate the effectiveness of our framework. First, we show that our method remains competitive with previous work [4, 6] on the standard Metaworld benchmark. Next, we augment the reach task of MetaWorld with obstacles, which have been never seen during pretraining, and compare performance against an augmented version of DP. We continue with real-world robot experiments to showcase our method’s robustness and generalizability. Finally, we elucidate our design choices by ablation studies.
@@ IV-A Replanning Frequency Evaluation on MetaWorld
In this section, we evaluate RA-DP on the full suite of MetaWorld tasks [29] to confirm its competitiveness to SOTA approaches despite achieving higher replanning frequencies. The tasks are categorized into four difficulty levels following [30]. As benchmarks are done in obstacle-free environments, all guidance signals are set to zero.
Experiments Setup. Expert demonstrations are obtained using the heuristic policies from MetaWorld for both state and point cloud observations. We collect only 10 demonstrations for all tasks following [6]. To ensure a fair comparison, we maintain the same resolution for images and depth maps as in the previous works. For state-based policies, we benchmark against Diffusion Policy (DP) [4], and for point cloud observations, we benchmark against 3D Diffusion Policy (DP3) [6].
@@
Results. The results are summarized in Table I. For both state-based and point cloud-based models, RA-DP outperforms SOTA methods with a large margin on success rate across most difficulties. The only exception is for Very Hard state-based tasks, which we attribute to insufficient training (3,000 epochs).
@@ IV-B Adaptation to Unseen Environmental Changes
Experiments setup. We collect 4,000 state-based demonstrations in an obstacle-free environment to train all models. In order to benchmark for collision avoidance, we modified the MetaWorld’s reach-v2 task. Specifically, a randomly placed spherical obstacle (static or dynamic) simulates an object to be avoided, such as a hand (see Fig. 3 for an example). We enhance DP [4] with training-free guidance integration as the baseline model, Guided-DP. Both RA-DP and Guided-DP use the same obstacle avoidance function to generate signals. In our experiments, we set $\lambda=50$, $\omega=100$ and $r=0.6$.
@@ IV-C Real Franka Arm Experiments and Demonstrations
Figure 4: Evaluation of RA-DPon real-world manipulation tasks where the robot interacts with human language instructions to avoid obstacles not been seen during training. A: Dynamic obstacle avoidance. B: Static obstacle avoidance. C1-C4: Static obstacle avoidance with human interference of stacking and removing different objects.
@@ IV-D Ablation Study
Influence of Dynamic Obstacle Speed. As a higher obstacle velocity presents greater challenges for collision avoidance. Here, we test the limit of RA-DP’s replanning ability with various obstacle speed on the dynamic obstacle avoidance task described in Section IV-B. We evaluate the success rate and the number of collision steps at various speeds, as shown in Fig. 6. We define collision steps as the number of simulation time steps the end-effector collided with the obstacle. A collision time step of zero Success. For clarity, the obstacle speed is normalized to a range of 0 to 1. We can observe that as the obstacle speed increases, the success rate declines, and the number of collision steps increases for the trajectories that reach the goal but do not fully avoid the obstacle. When the normalized speed exceeds 0.14, the replanning frequency of RA-DP becomes insufficient, causing the success rate to drop to zero.
@@
Mixed Noise Schedule. The mixed noise schedule controls the balance between monotonic and independent noise injection schedules during training, see Section III-B. A ratio of $0.0$ represents a purely monotonic schedule, while a ratio of $1.0$ denotes a purely independent schedule. Adjusting this ratio modulates the probability distribution between the two schemes.
We evaluate our method on a subset of MetaWorld’s easy tasks, using mixed ratios from $0.0$ to $1.0$ in increments of $0.2$. We found that while an independent noise schedule improved training convergence, the discrepancy between training and inference (which used a monotonic noise schedule) degrades test performance. Thus a mixture of two yields the best results. Empirically, a ratio of $0.6$ yields the best performance.
@@ V Conclusion
In this paper, we proposed a novel diffusion-based framework for training-free high-frequency replanning. Our method integrates a training-free guidance module, allowing generalization to novel dynamic feedbacks during inference. Additionally, an action queue mechanism is introduced to allow high-frequency replanning. Via empirical evaluation, we showed that the proposed approach achieves state-of-the-art success rate and replanning frequency in common and augmented benchmarks, and effectively handles complex dynamic interactions on a real robot. Although our study focused on collision avoidance, in theory, the proposed method can incorporate any type of training-free guidance signals. For instance, we can design differentiable functions that enforces a trajectory passage through specific waypoints or dynamically constrain the robot’s movement speed.

### v2

@@ Abstract
Diffusion models exhibit impressive scalability in robotic task learning, yet they struggle to adapt to novel, highly dynamic environments. This limitation primarily stems from their constrained replanning ability: they either operate at a low frequency due to a time-consuming iterative sampling process, or are unable to adapt to unforeseen feedback in case of rapid replanning. To address these challenges, we propose RA-DP, a novel diffusion policy framework with training-free high-frequency replanning ability that solves the above limitations by adapting to unforeseen dynamic environments. Specifically, our method integrates guidance signals, which are often easily obtained in the new environment during the diffusion sampling process, and utilizes a novel action queue mechanism to generate replanned actions at every denoising step without retraining, thus forming a complete training-free framework for robot motion adaptation in unseen environments. We conduct extensive evaluations in both common simulation benchmarks and real-world environments. Our results indicate that RA-DP outperforms the state-of-the-art diffusion-based methods in terms of replanning frequency and success rate. At the end, we show that our framework is theoretically compatible with any training-free guidance signal, hence increasing its applicability to a wide range of robotics tasks.
@@ I Introduction
Diffusion models have gained widespread attention for their effective performance in various tasks, such as image generation [1], video synthesis [2], and text-to-image conversion [3]. The ability of diffusion models to learn complex data distributions has made them a powerful tool in creative and predictive applications. More recently, diffusion models have been employed for learning robotic control tasks [4, 5, 6, 7]. In this context, they have demonstrated the expressiveness needed for complex manipulation tasks, stable training, and adaptation to higher-dimensional action spaces. These characteristics make diffusion models ideal for addressing a wide range of challenges in robotic control.
@@
Figure 1: Comparison between Diffusion Policy (DP) and RA-DP. DP has limited replanning frequency and fails to adapt to unforeseen dynamic observations, whereas RA-DP replans actions using the latest dynamic feedback at every denoising step in a training-free manner. The lighter the shade of blue, the lower the level of Gaussian noise. A white action represents an executable action.
For robot control, the authors of [4] propose Diffusion Policy (DP), which typically generates 16 actions by iterative sampling, executes the first 8 actions, and then replans conditioned on the new observations. DP has replanning frequency of $\approx 37$ Hz, which is substantially slower compared to standard low-level robotic control systems. As a result, DP can create a bottleneck for the entire system. One way to increase the replanning rate is through shortening of prediction and action horizons at the cost of lowering the success rate. Alternatively, one can resort to fast diffusion sampling techniques, such as consistency models [7], to improve frequency. However, these approaches still struggle to adapt to novel environment feedback, such as unforeseen obstacles.
@@
RA-DP establishes a fixed-length action queue. During training, each action within the queue is independently perturbed with varying noise levels, unlike previous methods that use uniform noise levels across actions. A standard U-Net is trained to predict clean actions conditioned on slowly changing feedback (e.g., a fixed goal position). During inference, RA-DP maintains an action queue with monotonically increasing noise levels. Other than conditioning on the foreseen, slow-changing feedback during the training stage, each action is also seamlessly steered by the training-free guidance from possible rapid changing environment feedback. RA-DP removes the executable clean action from the front of the queue (dequeue) and appends a new noisy one to the tail at every denoising step (enqueue), thereby achieving high-frequency training-free replanning. For example, in goal-reaching tasks in cluttered environments, RA-DP is able to incorporate live obstacle data to dynamically guide the actions.
@@
  • We propose RA-DP, a novel diffusion-based framework for robotic manipulation with a training-free high-frequency replanning to adapt to fast-changing environments.
@@ II-A Diffusion Models in Robotics
While these diffusion-based approaches excel in static or slow-changing environments, they fall short in rapidly evolving real-world scenarios. This is primarily due to their limited replanning frequency, hindering their application to tasks, such as collision avoidance in dynamic settings. To address this issue, we propose a novel high-frequency replanning diffusion-based framework that benefits from a training-free loss-based guidance in order to enhance adaptability and efficiency of the controller in highly dynamic environments.
@@ II-C Fast Replanning in Diffusion-based Controllers
Diffusion models [16, 17] generate high-quality samples through iterative refinement from Gaussian noise. This iterative process, however, leads to low replanning frequencies, limiting diffusion models real-time performance. To overcome this challenge, the approaches, such as DDIM [18], DPM-Solver [19], and Genie [20] use advanced ordinary differential equation (ODE) solvers to reduce the number of sampling steps needed for high-quality samples. Despite such an improvement, these methods fail to achieve the high-frequency replanning required for low-latency decision-making applications.
A more recent consistency policy [7] leverages diffusion model distillation to enable single-step sampling via direct noise-to-data mappings. Although this approach facilitates fast replanning, it has a number of shortcomings. The additional distillation stage applied to pretrained diffusion policies increases computational cost; The single-step sampling process compromises accuracy compared to iterative approaches; Lastly, the single-step architecture lacks intermediate latent representations, precluding the application of training-free, loss-based guidance techniques [10] necessary for adapting to unforeseen conditions. This is due to the fact that applying guidance directly to the data space either causes divergence or fails to influence the sampling process effectively.
In this work, we propose a significantly different approach, in which we integrate the action queue with different noise levels for high-frequency replanning. More specifically, our training and inference pipeline denoises the action queue with monotonically increasing noise levels, enabling single-step action generation without compromising sample quality. Additionally, we retain intermediate latent representations for training-free, loss-based guidance, thereby supporting adaptation to novel scenarios during inference. These characteristics make our proposed method compatible with any existing fast diffusion solvers.
@@
Similarly, hierarchical policies, such as Yell At Your Robot (YAY) [22], utilize high-level human language interventions as real-time corrective feedback for low-level policy errors. However, YAY’s VAE-based language-conditioned behavior cloning policy exhibits constrained action generation capabilities compared to diffusion-based approaches. Furthermore, while YAY requires continuous training to integrate language feedback, our method employs a training-free guidance mechanism.
@@ III Proposed Method
We begin by introducing the notation used throughout this paper and review the key aspects of Diffusion Policy (DP) in Section III-A. Then, in Section III-B, we present the training method of our RA-DP framework, which manages actions with independent noise levels during forward diffusion processes as opposed to how DP operates. Lastly, in Section III-C, we provide a detailed explanation of the inference process of our framework and demonstrate how it effectively achieves adaptive high-frequency replanning.
@@ III-A Notation and Diffusion Policy (DP)
We emphasize that the standard DDPM training used in diffusion treats the entire prediction sequence $A_{t}$ as a single example, where each action $a_{t}$ in the sequence is perturbed equally with the same noise level during the forward diffusion process. Instead, if we sample a sequence of actions during inference with varying noise levels, the policy trained with standard DDPM diverges and fails to predict normal actions. This is due to the mismatch between training and inference paradigms. However, we argue that varying noise level is needed to achieve real-time close-loop control.
@@ III-C Adaptive High-frequency Replanning Process
Finally, we update $\hat{A}_{t}^{\mathbf{k-1}}$ by appending a new noisy action $\hat{a}_{t+H}^{H\tau}$ to get the new action queue $\hat{A}_{t+1}^{\mathbf{k}}$. Leveraging our proposed action queue sampling method, at each time step $t$ after just one denoising step, we obtain a noiseless executable action $\hat{a}_{t}^{0}$ conditioned on the most recent observation $O_{t}$.
High-Frequency Replanning. Let $\Delta_{p}$ denote the cost of a one-step DDIM sampling and $\Delta_{a}$ the cost of a one-step action execution. Then, the replanning frequency of our framework is given by $\frac{1}{\Delta_{p}+\Delta_{a}}$. Notably, our method preserves sampling quality unlike other single-step diffusion samplers, as the total number of sampling steps for each action remains $H$.
@@
Adaptation to Novel Environment Feedback. Beyond conditioning on the slow-changing observation $O_{t}$, high-frequency replanning frameworks must swiftly respond to unforeseen environmental feedback. To this end, we propose to integrate a training-free loss-based guidance module [25, 26, 9], which outputs guidance gradients from arbitrary predefined constraints that influences the diffusion sampling process, thus achieving adaptation to novel environment feedback (see Fig.2). Moreover, the guidance module can be seamlessly integrated with the aforementioned action queue mechanism.
We describe the guided diffusion model from the perspective of score matching-based generative model. Here, both the forward and reverse diffusion processes are formulated as stochastic differential equations (SDEs), for which we train the U-Net to parameterize the score function $\nabla_{A^{\mathbf{k}}}\log p_{\theta}(A^{\mathbf{k}})$. However, a guided sampling process aims to sample $A^{\mathbf{k}}$ conditioned on the guidance signal $\mathcal{G}$, i.e., it requires the following score function:
@@ IV Experiments
We evaluate RA-DP using simulated and real-robot scenarios to highlight the importance of high-frequency replanning and demonstrate the effectiveness of our framework. First, we show that our method remains competitive with previous work [4, 6] on the standard Metaworld benchmark. Next, we augment the reach task of MetaWorld with obstacles, which have never been seen during pretraining, and compare our method to an augmented version of DP with guidance. We continue with real-world robot experiments to showcase our method’s robustness and generalizability. Finally, we highlight the impact of our design choices by conducting ablation studies.
@@ IV-A Replanning Frequency Evaluation on MetaWorld
In this section, we evaluate RA-DP on the full suite of MetaWorld tasks [29] to confirm its competitiveness to SOTA approaches while achieving higher replanning frequencies. The tasks are categorized into four difficulty levels following [30]. As benchmarks are done in obstacle-free environments, all guidance signals are set to zero.
Experiments Setup. Expert demonstrations are obtained using the heuristic policies from MetaWorld for both state and point cloud observations. We collect only 10 demonstrations for all tasks following [6]. To ensure a fair comparison, we maintain the same resolution for images and depth maps as in the previous works. For state-based policies, we benchmark against Diffusion Policy (DP) [4], and for point cloud observations, against 3D Diffusion Policy (DP3) [6].
@@
Results. The results are summarized in Table I. For both state-based and point cloud-based models, RA-DP outperforms the SOTA methods with a large margin on success rate across most difficulties. The only exception is for Very Hard state-based tasks, which we attribute to insufficient training (3,000 epochs).
@@ IV-B Adaptation to Unseen Environmental Changes
Experiments setup. We collect 4,000 state-based demonstrations in an obstacle-free environment to train all models. In order to benchmark for collision avoidance, we modified MetaWorld’s reach-v2 task. Specifically, a randomly placed spherical obstacle (static or dynamic) simulates an object to be avoided, such as a hand (see Fig. 3 for an example). We augment DP [4] with training-free guidance integration as the baseline model, Guided-DP. Both RA-DP and Guided-DP use the same obstacle avoidance function to generate signals. In our experiments, we set $\lambda=50$, $\omega=100$ and $r=0.6$.
@@ IV-C Real Franka Arm Experiments and Demonstrations
Figure 4: Evaluation of RA-DP on real-world manipulation tasks where the robot interacts with obstacles not seen during training. A: Dynamic obstacle avoidance. B: Static obstacle avoidance. C1-C4: Static obstacle avoidance with human interference of stacking and removing different objects.
@@ IV-D Ablation Study
Influence of Dynamic Obstacle Speed. A higher obstacle velocity presents greater challenges for collision avoidance. Here, we test the limits of RA-DP’s replanning ability with various obstacle speeds on the dynamic obstacle avoidance task described in Section IV-B. We evaluate the success rate and the number of collision steps at various speeds, as shown in Fig. 6. We define collision steps as the number of simulation time steps that the end-effector collided with the obstacle. A collision time step of zero is Success. For clarity, the obstacle speed is normalized to a range of 0 to 1. We can observe that as the obstacle speed increases, the success rate declines, and the number of collision steps increases for the trajectories that reach the goal but do not fully avoid the obstacle. When the normalized speed exceeds 0.14, the replanning frequency of RA-DP becomes insufficient, causing the success rate to drop to zero.
@@
Mixed Noise Schedule. The mixed noise schedule controls the balance between monotonic and independent noise injection schedules during training ( see Section III-B). A ratio of $0.0$ represents a purely monotonic schedule, while a ratio of $1.0$ denotes a purely independent schedule. Adjusting this ratio modulates the probability distribution between the two schemes.
@@
We evaluate our method on a subset of MetaWorld’s easy tasks, using mixed ratios from $0.0$ to $1.0$ in increments of $0.2$ as shown in Table III. We can see that while an independent noise schedule improves training convergence, the discrepancy between training and inference (which used a monotonic noise schedule) degrades test performance. Thus a mixture of two yields the best results. Empirically, a ratio of $0.6$ yields the best performance.
@@ V Conclusion
In this paper, we proposed a novel diffusion-based framework for training-free high-frequency replanning. Our method integrates a training-free guidance module, allowing generalization to novel dynamic feedbacks during inference. Moreover, our method benefits from an action queuing mechanism, which allows high-frequency replanning. Via empirical evaluation in common and augmented benchmarks as well as real-world scenarios, we showed that the proposed approach achieves state-of-the-art performance while maintaining high replanning frequency, and can effectively handle complex dynamic interactions. Although our study focused on collision avoidance, in theory, the proposed method can incorporate any types of training-free guidance signals. For instance, we can design differentiable functions that enforce a trajectory passage through specific waypoints or dynamically constrain the robot’s movement speed.